A bulet of mass m and velocity v strikes a mass M connected to a spring (k=spring constant) [the spring is placed horizontaly and other end is fixed.]and get embeded to it . Find the amplitude of oscilation.

Grade:12

2 Answers

vikas askiitian expert
13 years ago

first applying conservation of momentam

 mv = (M+m)Vf

 Vf = mv/m+M  

now , kinetic energy of block + mass = KE

   KE = (m+M)Vf2 /2

  this system will stop moving when its all KE is converted into PE

 so , KE = PE

      KX2 = (M+m)Vf2

       X =[ m2v2 /(M+m)K ]1/2  = max amplitude

vikas askiitian expert
13 years ago

please approve my answers if they are correct